OpenStudy (anonymous):

Just checking my answer: Solve the equation: (3x - 2)(7x + 1) = 0. A. {2/3, -1/7} B. {-2/3, 1/7} C. 2/3 only D. -1/7 only E. no solution

OpenStudy (anonymous):

I think it's A

sam (.sam.):

(3x-2)(7x+1)=0 (3x-2)=0 (7x+1)=0 ----------------------------- 3x-2=0 3x=2 3x/3=2/3 x=2/3 ----------------------------- 7x+1=0 7x=-1 7x/7=-1/7 x=-1/7 ==================== x=2/3 , -1/7

OpenStudy (anonymous):

thnk u

OpenStudy (anonymous):

its A.